Leg 173 Operations Synopsis

To Site 1068

SITE 1067
(Proposed Site Iberia-9A)

The vibration-isolated television (VIT) camera was left deployed while the ship was moved from Site 1066 to Site 1067. No subsea cables were visible on the seafloor. Hole 1067A was spudded at 0100 hr 2 May 1997. It was at this time that the core tech noticed a spooling problem with the coax winch. While the ship was offset to the southwest in search of deeper water, an additional 30 m of line was let out and the problem was corrected. After about 2.5 hr, the ship was back on location and ready to re-spud the hole.

Hole 1067A was "officially" spudded at 0330 hr 2 May 1997. TheVIT was retrieved during a jet in test to a depth of 5092.8 mbrf or 60.8 mbsf. Drilling with an RCB/center bit commenced at 0600 hrs; however, 2.5 hr later, at a depth of 109.0 mbsf, the WKM valve developed a severe leak causing the drilling operation to be halted for less than 2 hr. Drilling continued to a depth of 5680.0 mbrf averaging 14.5 m/hr. Drilling weight on bit was limited to 15,000 lb to prevent over loading/torquing the core barrel latch and latch sleeve. A total of 44.75 hr were required to drill the hole down to 648.0 mbsf, including time to recover and inspect the center bit four times.

Continuous RCB coring began at this depth and, after encountering basement at approximately 764.0 mbsf, coring continued to 855.6 mbsf. Between Cores 17R and 18R, a short trip of the drill string was required to replace 10 stands of drill pipe fitted with aluminum pipe protectors with standard 5-1/2-in pipe. Given the length of pipe that was removed, this short trip raised the bit to level above that of the sediment-basement contact. A wash barrel was used while running the pipe back to the bottom of the hole, and Core 18R was taken using this same core barrel. Thus the numerous claystone and siltstone pebbles at the top of Core 18R are almost certainly derived from sediments overlying the sediment-basement contact.

When coring had proceeded to 855.6 mbsf, the original C-4 core bit had 101.6 rotating hours and the rate of penetration for Core 23R was down to 0.9 m/hr. A sepiolite mud pill was circulated and the drill string was pulled to a depth of 185.0 mbsf. During the trip an overpull of 50,000 lb was seen at a depth of 744.0 mbsf; however, this was the only tight spot in the hole. After rigging and dropping a free fall funnel (FFF), theVIT was deployed. On the first attempt at pulling clear of the seafloor the C-4 core bit snagged the 13-3/8 in. casing joint on the FFF. The bit was immediately lowered back down hole ~9 m and another attempt was made. This time the bit came clear of the FFF without incident and at 0600 hr 9 May 1997 the bit cleared the seafloor. The VIT was retrieved as the pipe trip continued to the surface and the bit was on deck at 1400 hr.

A new C-7 bit, and a rebuilt mechanical bit release (MBR), were being made up to the outer core barrel when a vertical crack ~2 in long was found that ran downward from the shoulder of the MBR box connection. The cracked MBR top connector, used previously in Hole 1065A, was replaced and the drill string was tripped to the seafloor. With the VIT deployed, the vessel maneuvered for 3.5 hr before reaching a position over the FFF. Reentry was made at 0415 hr on 10 May 1997. At the time of reentry, vessel heave was causing the weight indicator to fluctuate 20,000-40,000 lb. Once inside the hole, the drill string was lowered slowly until 0430 hr, when the bit reached a depth of 80.0 mbsf. At this point 20,000-25,000 lb weight on bit was taken.

The drill string could not be rotated or run in at normal speeds until the VIT was retrieved. Use of the heave compensator or rapid downward movement of the drill string could have jeopardized the safe recovery of the VIT system because the cable can wrap around the drill pipe. While retrieving the VIT (~45 min), the string was worked up and down slowly three times with 20,000-25,000 lb weight in an attempt to pass the obstruction and place the BHA in a less vulnerable position. When the drill string was picked up the second time, an ~20,000-lb loss in string weight was noted, and the drill string was pulled back to 5102 mbrf where the circulating head was made up and circulation started. The drill string was not moved again until the camera frame was recovered.

After recovering the VIT frame, a rotary core barrel was run in on the wireline. If the BHA had been intact, the core barrel should have landed at 5108 mbrf. Instead, the barrel repeatedly landed at 5032 mbrf or 76 m above the landing point. The core barrel was pulled back ~20 m and landed two more times at the same point. This indicated that there had indeed been a failure in the BHA, most likely in a drill collar connection. The wireline was retrieved, the top drive set back, and the drill string was pulled back to the surface. While pulling the pipe, the beacon was released and subsequently recovered at 1745 hr on 10 May 1997. The BHA cleared the rotary table at 1815 hr that same day officially ending Hole 1067A.

A BHA failure was confirmed in the connection between 8-1/4-in drill collars #7 and #8. While the vessel was moved in DP mode to Site 1068 (Iberia-9C), the two remaining 8-1/4-in. drill collars and the tapered drill collar were inspected using magnetic particle techniques with no positive indications of any cracks. As a result of the BHA failure, Hole 1067A was abandoned short of the primary depth objective.

Back to Top

SITE 1068

After finishing at Site 1067, the ship was moved in DP mode to alternate Site 1068 (Iberia-9C; not in original prospectus) located ~1600 m to the west of Hole 1067A. Once on location, the VIT camera system was deployed to greater than 5000 m water depth to detorque the coaxial cable. During this time, the new drill collars for the BHA were picked up and the RCB core barrels were spaced out. A beacon was deployed at 1845 hr on 10 May 1997; however, the beacon signal became erratic. At 2100 hr that same day, a second beacon was deployed.

Because of spooling problems with the coax winch, the TV system was near the seafloor for about 9 hr while we tried to correct gaps in the spooling of the cable on the winch. No indication was observed of subsea cables during the survey.

The drill string was tripped to bottom and Hole 1068A was spudded at 1430 hr 11 May 1997. A rig-floor corrected seafloor depth of 5055.0 m was established. Drilling ahead with the RCB center bit assembly continued while checking the center bit approximately every 100 m, beginning at 300 mbsf. Use of the 5-1/2-in. drill pipe with the aluminum wear knots was begun at a depth of ~600 mbsf. Drilling proceeded until 1630 hr on 14 May 1997, when a depth of 711.3 mbsf was reached. Originally the plan had been to drill to a depth of 780.0 mbsf before initiating coring; however, because of the limit to how much weight can be placed on the bit while drilling with a center bit, the increasingly slow penetration rate (2.8 m/hr) meant that coring would take equal or less time than drilling. Therefore, the center bit was recovered, and a core barrel was deployed. Coring with the RCB continued at an average ROP of 3.6 m/hr until basement was reached in Core 15R at an approximate depth of 852.0 mbsf. Coring continued into the basement rocks including a several-meter-thick band of dark bluish black serpentinite breccia that was extremely difficult to core. This material packed very easily, jamming the core catcher, and slowing the ROP to 1.0-2.0 m/hr. With depth, the formation graded into massive serpentinized peridotite. Coring was eventually suspended after Core 29R at a depth of 955.8 mbsf because of time constraints and the uncertainty about how much deeper the hole would have to progress before unaltered basement rock (peridotite) could be reached.

During the wiper trip in preparation for wireline logging, the driller noted constant 25,000 lb drag until reaching a depth of 781.0 mbsf. During the trip, the top drive had to be picked up at a depth of 775.0 mbsf. In general, high torque was noted from 899.0 mbsf to 915.0 mbsf, and there were three exceptionally tight spots that required heavy reaming and high torque at depths of 775.0 mbsf, 899.0 mbsf, and 908.0 mbsf. There was 6.0 m of fill on bottom. In preparation for logging, the hole was flushed with 30 bbl of sepiolite mud, the bit was released, the sleeve shifted, and the hole was displaced with another 265 bbl of sepiolite mud. The drill pipe was pulled to a logging depth of 114.0 mbsf and at 0415 hr on 21 May 1997, the crew began rigging up for logging. Because the hole was in poor condition, particularly in the basement rocks, and because of the heavy reaming required during the wiper trip, the side entry sub was not used.

Wireline logging proceeded with the first suite of logging tools, the triple combo, consisting of NGS/DIT/HLDT/APS and the Lamont temperature tool. The first run had trouble passing a tight spot in the hole at 520.0 mbsf. After unsuccessfully expending some effort to pass this point, the hole was logged back up to the open ended pipe (114.0 mbsf). After recovering the first suite of tools, the pipe was lowered across the bridge placing the open ended pipe at a depth of 614.0 mbsf. This was the deepest the pipe that could be placed without severely jeopardizing the pipe string. The triple combo logging suite was again run in the hole, this time reaching a depth of 770.0 mbsf before again reaching an obstruction. The hole was logged up from that point, the tools recovered, and further wire line logging efforts were abandoned. Upon concluding wireline logging operations, the pipe was pulled clear of the seafloor at 2200 hr on 21 May 1997. A pipe overpull of 25,000 lb was noted on the first two stands of drill pipe. After pulling clear of the seafloor, the first beacon was commanded to release but did not surface. Once the second beacon was released and recovered aboard at 2400 hr, the vessel began to slowly DP move to alternate Site 1069. The pipe trip continued during the DP move. The mechanical bit release reached the rig floor at 0530 hr on 22 May 1997, officially ending Hole 1068A.

Back to Top

To 173 Operations: Sites 1069 and 1070

To 173 Table of Contents

ODP Publications

ODP Home